OpenStudy (anonymous):

Mass, in grams, of 0.185mol sucrose (C12H22O11).

OpenStudy (aaronq):

given the formula: n=m/M (i.e. moles=mass/molecular mass) i'm sure you can solve it

OpenStudy (anonymous):

no i cant

OpenStudy (aaronq):

do you know how to solve algebraic problems? eg 4=x/6

OpenStudy (aaronq):

?

OpenStudy (anonymous):

Multiply the number of each element (such as 12 carbons) by it's atomic mass and add up all the masses to get the molar mass of sucrose. Then with that, do dimensional analysis

OpenStudy (anonymous):



OpenStudy (aaronq):

@Doublecook dude, don't tell people the answers straight up. Let them work it out, unless it seems to difficult. You're contributing to the degradation of education